OpenStudy (anonymous):

Is anyone here the master of work and energy and can help a freshman out? Regarding spins

OpenStudy (anonymous):

There's a 2kg mass moving on a surface which in the end there's a spin. The distance between A-B is 5 meters and ONLY between A-B there's friction (meaning under the spin there's no friction) What should be the initial velocity of the mass so it will come back to point A? the coefficient of friction is 0.2.

OpenStudy (anonymous):

I shall cherish you all whoever is capable of solving it!!!!

OpenStudy (anonymous):

I assume the spring is fixed in place and the 2 Kg mass compresses it and the spring stopping it will reverse its motion and the mass leaves the spring with the same speed as it engaged it. The initial KE energy must equal the work done in overcoming friction in both directions.

OpenStudy (michele_laino):

I think that the initial kinetic energy of your mass has to be greater than the work by the friction force acting in both direction, namely: \[\frac{ 1 }{ 2 }mv ^{2}>2(\mu*m*g*d)\] where \[\mu \] is the coefficient of sliding friction

OpenStudy (michele_laino):

of course I suppose that the collision between the mass and the spring is elastic

OpenStudy (anonymous):

Thank you! but what the the answer then?

OpenStudy (anonymous):

Approximately 6.26 m/s

OpenStudy (anonymous):

you're brilliant!!! Can you please show me the solution?

OpenStudy (anonymous):

So I will be able to do it myself for next time

OpenStudy (anonymous):

The problem with assuming that the kinetic energy must be greater than the work done to over come friction in both directions is that the mass would land at a position beyond A. Set the equation for kinetic energy equal to 2 * the equation for the friction force: \[1/2 *mv^2 = 2(\mu m g d)\]

OpenStudy (anonymous):

plug in everything: 1/2*(2kg)v^2 = 2[(0.2)(2kg)(9.81m/s^2)(5m)]

OpenStudy (anonymous):

then solve for v

OpenStudy (anonymous):

You are truly my hero!!!!! Thank you so much!!!!!!!!!!!!!!!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Can I ask you something short?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Based on the K of the spin is 52.3N, what would be the Delta X of the spin when the box hits it? Which equation should I use?

OpenStudy (anonymous):

when you say spin, are you referring to a spring?

OpenStudy (anonymous):

0.skx^2=W?

OpenStudy (anonymous):

spring yes sorry

OpenStudy (anonymous):

so if the spring constant, k, is 52.3 N/m, are you asking what the force is when the box hits it?

OpenStudy (anonymous):

what would be the constriction of the spring when the box hits it

OpenStudy (anonymous):

*hits the spring

OpenStudy (anonymous):

Use hooke's law and solve for X

OpenStudy (anonymous):

I tried to do it and got confused...

OpenStudy (anonymous):

KX^2*0.5=W

OpenStudy (anonymous):

Am I correct? And to work it will be coefficient of friction*mg*5*cos180?

OpenStudy (anonymous):

are you talking about the elastic potential energy equation?

OpenStudy (anonymous):

I thought that would fit the best here, am I completely wrong?

OpenStudy (anonymous):

set that equal to kx and solve for x

OpenStudy (anonymous):

0.5KX^2-0.5MV^2=coefficient of friction*mg*cos180*5 meters?

OpenStudy (anonymous):

Is this the calculation I should do?

OpenStudy (anonymous):

BTW thank you so much for everything

OpenStudy (anonymous):

Am I right this time?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think I've got it. Thanks allot!!!!! You've truly made my day brighter!

OpenStudy (anonymous):

But can you just approve my equation?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Your equation is correct and can be written as \[\frac{ 1 }{ 2 }kx ^{2} =\frac{ 1 }{ 2 }mv ^{2}-5\mu mg\]where v is the initial velocity of the box, x is the net compression of the spring. The equation states that the energy stored in the spring is equal to the net energy available just before the spring and box engage which is also equal to the beginning available energy (KE of the Box) minus the work done before engaging the spring. Is that clear?